# An Information-Theoretic Approach to Understanding Transformers’ In-Context Learning of Variable-Order Markov Chains

### 28 Mar 2026 v2

@@ 1 INTRODUCTION
The CTW algorithm needs the counts for the suffixes, and the higher layers of our transformer construction are mostly performing information aggregation. We conjecture that 2-layer transformers can perform reasonably well because even raw counts or partially aggregated information would allow a close-to-optimal approximation. We then consider such a simplified 2-layer transformer, by providing one feed-forward (FF) layer with the probability estimates and various configurations of the counts directly. The FF layer can be trained to approximate the proper blending coefficients using this information. We implement several synthetic transformer layers, and experiments show that their performances indeed match our conjectures.
@@
    Our construction allows us to investigate the relative performance insensitivity to the number of layers, i.e., why $2$-layer transformers perform well.
@@ 3 IN-CONTEXT LEARNING OF VOMCS
We train transformers to conduct experiments. A ternary alphabet $|\mathcal{A}|=3$ is adopted, and transformers of context window size $N$ are pretrained on data sequences of length-$N$ generated using CTs randomly sampled from the given prior $\pi_{\text{CTW}}$ parameterized by $\bm{\alpha}=0.5$, $\lambda=0.15$ and a maximum tree depth $D$; training details are given in Appendix D. The training loss is the sum of canonical cross-entropy losses in the context window, or equivalently, the sum of data compression rates. During testing, given a source sequence of length-$N$ generated from an unknown VOMC with an order at most $D$, can the transformer learn this sequence efficiently, i.e., at a loss close to the optimal value over the whole context window? In Fig. 3, we show the ICL performances (the compression rates, equivalent to the cross-entropy losses) of the trained transformers with various numbers of layers, and the CTW and KN algorithms for $N=1536$ and CT maximum order $D=5$. The transformers have 8 attention heads with embedding dimension $E=128$. The horizontal axis is the position in the context window, and the rates indicate how well transformers and other algorithms learn the underlying VOMCs in context.
@@
Previous works Edelman et al., (2024); Nichani et al., (2024) showed that a two-layer attention-only network (i.e., without attention layer normalization or the FF layer) can effectively perform ICL for FOMCs, where the second layer only requires a single attention head. Moreover, the authors of Rajaraman et al., 2024a empirically observed that two-layer single-head transformers can perform ICL-FOMCs of at least order 4, and constructed a three-layer single-head attention-only network (without an FF layer but with attention layer normalization) that can perform ICL-FOMCs. We conducted experiments to verify whether these continue to hold for ICL-VOMCs, and the results are shown in Fig. 4, where the bracket after TF (transformer) indicates the number of heads in each layer. Observe that attention-only networks perform considerably worse than CTW and a full four-layer transformer. Two-layer single-head transformers, and three-layer single-head transformers without the FF layer, also perform poorly, particularly the former. These results show that the mechanisms given in Edelman et al., (2024); Nichani et al., (2024); Rajaraman et al., 2024a for ICL-FOMCs cannot explain well how transformers perform ICL-VOMC, and the feedforward networks need to play a more significant role.
@@ 4.2 The Bayesian CTW Algorithm
    where $qs$ is the string by appending symbol $q\in\mathcal{A}$ before the suffix $s$. This recursion weights $p^{e}$ at a node with $p^{w}$’s at its children nodes.
@@ 4.3 Analysis of Attention Maps
We analyzed the attention maps of the trained transformers, where a few distinguished patterns emerge. One pattern is solely relative-position dependent. In the first two panels of Fig. 5 (top), we observe off-diagonal stripes for these two attention heads, which are a few positions below the main diagonal. They can be a single off-diagonal or a collection of several off-diagonals. This indicates that the query position is attending positions at a few fixed but close distances ahead of itself. This pattern usually appears in the first or second layers of the transformers. Combining with the suffix structure in compression algorithms such as CTW, such an attention pattern suggests the suffix is being copied into the current query position for subsequent processing. The off-diagonal stripes may have a width greater than $1$, as shown in the second panel, which can be viewed as copying a mixture of the tokens in the suffix, suggesting the flexibility of transformers in forming certain “soft" suffixes.
Another pattern, shown in the third panel, has more sophisticated spotty patterns, and the attention appears to depend more explicitly on the current token features instead of the position alone, and they usually appear in the second layer or above in the transformers. Taking query positions 350 and 362 for the attention head shown in the third panel of Fig. 5 (top), we plot in Fig. 5 (bottom) the positions in the data sequence that match their suffixes of length-$3$ using the stem plots with a black circle on top, and the attention values as the red stems with the diamonds on top. The positions match perfectly, though the attention weights have some variations among them. This attention pattern suggests that it is collecting statistical information for those positions with the matched suffix. Several attention heads present similar patterns but with different suffix lengths.
@@ 4.4 Interpretation via Constructions
We next provide a construction of $(2+D)$-layer transformer with sufficient representation power in the FF layer that can essentially approximate CTW, which demonstrates the capacity of transformers. The first two layers are motivated by the attention map patterns observed in Section 4.3, which collect statistics needed for Theorem 2; the last $D$ layers are induction layers imitating the CTW procedure.
@@
It is important to note that for ICL of FOMCs, two layers that collect forward statistics $\vec{g}_{i,M^{\prime}}$ with $M^{\prime}=D+1$ are sufficient (Edelman et al.,, 2024). However, for the ICL-VOMC task, the underlying CT structure is unknown; therefore, collecting such simple statistics is no longer sufficient, and it is also the reason that the FF layer is important for VOMCs while less so for FOMCs. As indicated in Theorem 2, the information of counting statistics $\vec{n}_{i,s_{i,l}}$ is important since the weights heavily depend on $\vec{n}_{i,s}(a)$. Yet due to the softmax function, only (normalized) probability values can be obtained instead of the counts. With the backward statistics $\vec{g}^{\leftarrow}_{i,s}$, $\vec{n}_{i,s_{i,l}}$ can be derived as
@@
Taking $M=M^{\prime}-1=D$, after the statistics collection layer, a sufficiently wide FF layer with ReLU activation can compute (by the universal approximation capability) (Cybenko,, 1989; Hornik et al.,, 1989).
@@
There exists an $A$-head transformer
layer that can perform the induction: Takes $\vec{H}^{(\ell)}$ in (7) as input and outputs $\vec{H}^{(\ell+1)}$, and the final output layer taking $\vec{H}^{(D+3)}$ as input can output the $A$-dim Bayesian optimal vector $\sum_{l=0,\ldots,D}~\omega_{n,l}\vec{p}_{n,s_{n,l}}.$
@@
The construction is certainly not unique. Consider a 2-layer transformer with a large embedding space, where the number of attention heads scales like $|\mathcal{A}|^{D}$. Each attention head could be precoded to collect the statistics for each possible suffix of length $\leq D$, and all such information can be stored in $\vec{h}_{i}^{(3)}$, assuming a sufficiently large embedding space (or a more delicate embedding structure in a smaller space). Each $\vec{h}_{i}^{(3)}$ would then have all the information on the context tree, on which a wide FF layer could perform the full CTW algorithm to produce the next token distribution. We conscientiously forgo constructions like this, because in large models with $\mathcal{A}$ in the tens of thousands and $D$ in the tens or hundreds, scaling as $|\mathcal{A}|^{D}$ is unrealistic, and the computation of the FF layer would be extremely complex and hard to interpret. More discussions on potential tradeoffs between layers and attention heads can be found in Rajaraman et al., 2024a .
@@ 4.5 Interpreting Two-Layer Transformers
We next consider how two-layer transformers can learn VOMCs in context. Our construction given above requires $D+2$ layers, and we conjecture that the statistical information collected in the second layer can be used to approximate the optimal weights $\omega_{n,l}$ with an FF network.
To verify our conjecture, we implement the first two constructed transformer layers but control the statistics passed onto the FF layer. In Fig. 8, “TF 0-2” denotes the canonical 2-layer transformer; “TF 2-0” is the transformer with two constructed layers with output $\vec{a}_{i}^{(2)}$ in (6), followed by a trainable FF layer (the FF layer in the second layer of the transformer) and an output layer. The version “TF 2-0 w/o counts” does not contain $\vec{g}_{i-1,M^{\prime}}^{\leftarrow}$ or $\vec{pos}_{i}$ in $\vec{a}_{i}^{(2)}$; the version “TF 2-0 total counts only” does not contain $\vec{g}_{i-1,M^{\prime}}^{\leftarrow}$ in $\vec{a}_{i}^{(2)}$ and $\vec{pos}_{i}$ is replaced by the total count $i$; “TF 2-0 w/ all counts” replaces $\vec{g}_{i-1,M^{\prime}}^{\leftarrow}$ and $\vec{pos}_{i}$ with $\{\vec{n}_{n,s_{n,l}}\}_{l=0}^{D}$ and $i$. Even though their performances are rather clustered, we can make the following observations: 1) The performances degrade as more counting information is removed from the representation, and the counting information is clearly important, 2) The performances of “TF 2-0” and “TF 2-0 w/ all counts” almost match exactly, indicating the main function of the backward statistics $\vec{g}^{\leftarrow}_{i-1,M^{\prime}}$ is to extract the counts, and 3) The performance of the canonical 2-layer transformer is similar to that of the constructed “TF 2-0” and “TF 2-0: w/ all counts” than those with less counting information. Therefore, the performances of these different implementations confirm the conjecture. Interestingly, Akyürek et al., (2024) proposed a similar learned-weight 22&33-gram algorithm to interpret how transformers learn finite automata, and suggested it may be optimal. In contrast, we obtain the (suboptimal) architecture by approximating the provable optimal transformer architecture.
@@ Appendix D Training Details
We choose the alphabet size to be $|\mathcal{A}|=3$ in the experiments. For training, we randomly generate $K=20000$ CTs of various depths (maximum order $D\leq 5$), and then for each CT leaf, we generate a probability distribution. Two different ways of generating these probability distributions are taken: the first approach is to use the Dirichlet distribution to sample such distributions, and the second approach is to randomly select some of the elements in the alphabet to have probability zero, and the others with i.i.d. random values before normalization. Different values of the Dirichlet parameter are tested but only the results do not appear to be sensitive to the choice. For each CT, a source sequence of a certain length (e.g., $N_{k}=5120$) is produced. The context window $N$ can vary, but in most cases, we set it at $512$ (except when $D=5$, we set it to be $1536$ to allow sufficient data collection in context). Each source sequence is segmented into $\lfloor N_{k}/N\rfloor$ training sequences.
@@ G.1 A New Formula for Bayesian Next Token Prediction
Discussion. Note that $p^{e}_{n,s},p^{w}_{n,s}$ can be efficiently calculated by the CTW procedure, and compared to calculate $\frac{P_{\pi_{\text{CTW}}}(x_{1}^{n+1}|x_{1-D}^{0})}{P_{\pi_{\text{CTW}}}(x_{1}^{n}|x_{1-D}^{0})}$ for each $x_{n+1}$ the extra computation besides the CTW procedure is $A$ times larger than that by Eq (10). The weighted average formula in Eq (10) gives a natural interpretation for the Bayesian optimal next token predicted probability. Each suffix along the root the leaf path $s_{n,0}-s_{n,1}-\cdots-s_{n,D}$ can potentially be the true suffix, i.e., $s_{n,l}\in\mathcal{L}(T)$, and $\vec{p}_{n,s_{n,l}}$ is in fact the Bayesian optimal next token prediction given $s_{n,l}\in\mathcal{L}(T)$.
The blending weights $\omega_{n,l}$’s are based on stopping probability $\lambda$, the information in the potential suffix path such as $p^{e}_{s_{n,s_{n,l}}}$ as well as the information from their siblings $p^{w}_{n,qs_{n,l-1}}$. We can interpret $p^{e}_{n,s}$ as the evidence (unnormalized likelihood) that $s\in\mathcal{L}(T)$, and $p^{w}_{n,s}$ as the evidence that $s\in T$, i.e., the underlying tree covers node $s$. Theorem 6 indicates that more weights are assigned to $s_{n,l}$ than $s_{n,l-1}$, i.e., $\omega_{n,l}>\omega_{n,l-1}$, if $\lambda$ is smaller (i.e., node $s_{n,l-1}$ is more likely to branch and thus less likely to be a leaf node), $p^{e}_{n,s_{n,l}}-p^{e}_{n,s_{n,l-1}}$ is larger (i.e., $s_{n,l}$ has more evidence than $s_{n,l-1}$) and $\sum_{q\in\mathcal{A}}\ell^{w}_{n,qs_{n,l-1}}-\ell^{w}_{n,s_{n,l}}$ is larger (i.e., $_{n,l}’s s$iblings have more evidence to explain the data and thus $s_{n,l-1}$ is less likely to be a leaf node). The indicator function $\mathbb{I}_{l=D}$ is due to the maximum depth constraint on the branching process. Nodes at level $l=D$ automatically stop the branching, i.e., the branching-stopping probability is $1$ for such nodes.

### 31 Mar 2026 v3

@@ 1 INTRODUCTION
The CTW algorithm needs the counts for the suffixes, and the higher layers of our transformer construction are mostly performing information aggregation. We conjecture that 2-layer transformers can perform reasonably well because even raw counts or partially aggregated information would allow a close-to-optimal approximation. We then consider such a simplified 2-layer transformer, by providing one feed-forward ($F$) layer with the probability estimates and various configurations of the counts directly. The $F$ layer can be trained to approximate the proper blending coefficients using this information. We implement several synthetic transformer layers, and experiments show that their performances indeed match our conjectures.
@@
    Our construction allows us to investigate the relative performance insensitivity to the number of layers, i.e., why 22-layer transformers perform well.
@@ 3 IN-CONTEXT LEARNING OF VOMCS
We train transformers to conduct experiments. A ternary alphabet $|\mathcal{A}|=3$ is adopted, and transformers of context window size NN are pretrained on data sequences of length-NN generated using CTs randomly sampled from the given prior $\pi_{\text{CTW}}$ parameterized by $\bm{\alpha}=0.5$, $\lambda=0.15$ and a maximum tree depth $D$; training details are given in Appendix D. The training loss is the sum of canonical cross-entropy losses in the context window, or equivalently, the sum of data compression rates. During testing, given a source sequence of length-NN generated from an unknown VOMC with an order at most $D$, can the transformer learn this sequence efficiently, i.e., at a loss close to the optimal value over the whole context window? In Fig. 3, we show the ICL performances (the compression rates, equivalent to the cross-entropy losses) of the trained transformers with various numbers of layers, and the CTW and KN algorithms for $N=1536$ and CT maximum order $D=5$. The transformers have 8 attention heads with embedding dimension $E=128$. The horizontal axis is the position in the context window, and the rates indicate how well transformers and other algorithms learn the underlying VOMCs in context.
@@
Previous works Edelman et al., (2024); Nichani et al., (2024) showed that a two-layer attention-only network (i.e., without attention layer normalization or the $F$ layer) can effectively perform ICL for FOMCs, where the second layer only requires a single attention head. Moreover, the authors of Rajaraman et al., 2024a empirically observed that two-layer single-head transformers can perform ICL-FOMCs of at least order 4, and constructed a three-layer single-head attention-only network (without an $F$ layer but with attention layer normalization) that can perform ICL-FOMCs. We conducted experiments to verify whether these continue to hold for ICL-VOMCs, and the results are shown in Fig. 4, where the bracket after TF (transformer) indicates the number of heads in each layer. Observe that attention-only networks perform considerably worse than CTW and a full four-layer transformer. Two-layer single-head transformers, and three-layer single-head transformers without the $F$ layer, also perform poorly, particularly the former. These results show that the mechanisms given in Edelman et al., (2024); Nichani et al., (2024); Rajaraman et al., 2024a for ICL-FOMCs cannot explain well how transformers perform ICL-VOMC, and the feedforward networks need to play a more significant role.
@@ 4.2 The Bayesian CTW Algorithm
    where qsqs is the string by appending symbol $q\in\mathcal{A}$ before the suffix $s$. This recursion weights $p^{e}$ at a node with $p^{w}$’s at its children nodes.
@@ 4.3 Analysis of Attention Maps
We analyzed the attention maps of the trained transformers, where a few distinguished patterns emerge. One pattern is solely relative-position dependent. In the first two panels of Fig. 5 (top), we observe off-diagonal stripes for these two attention heads, which are a few positions below the main diagonal. They can be a single off-diagonal or a collection of several off-diagonals. This indicates that the query position is attending positions at a few fixed but close distances ahead of itself. This pattern usually appears in the first or second layers of the transformers. Combining with the suffix structure in compression algorithms such as CTW, such an attention pattern suggests the suffix is being copied into the current query position for subsequent processing. The off-diagonal stripes may have a width greater than 11, as shown in the second panel, which can be viewed as copying a mixture of the tokens in the suffix, suggesting the flexibility of transformers in forming certain “soft" suffixes.
Another pattern, shown in the third panel, has more sophisticated spotty patterns, and the attention appears to depend more explicitly on the current token features instead of the position alone, and they usually appear in the second layer or above in the transformers. Taking query positions 350 and 362 for the attention head shown in the third panel of Fig. 5 (top), we plot in Fig. 5 (bottom) the positions in the data sequence that match their suffixes of length-33 using the stem plots with a black circle on top, and the attention values as the red stems with the diamonds on top. The positions match perfectly, though the attention weights have some variations among them. This attention pattern suggests that it is collecting statistical information for those positions with the matched suffix. Several attention heads present similar patterns but with different suffix lengths.
@@ 4.4 Interpretation via Constructions
We next provide a construction of $(2+D)$-layer transformer with sufficient representation power in the $F$ layer that can essentially approximate CTW, which demonstrates the capacity of transformers. The first two layers are motivated by the attention map patterns observed in Section 4.3, which collect statistics needed for Theorem 2; the last $D$ layers are induction layers imitating the CTW procedure.
@@
It is important to note that for ICL of FOMCs, two layers that collect forward statistics $\vec{g}_{i,M^{\prime}}$ with $M^{\prime}=D+1$ are sufficient (Edelman et al.,, 2024). However, for the ICL-VOMC task, the underlying CT structure is unknown; therefore, collecting such simple statistics is no longer sufficient, and it is also the reason that the $F$ layer is important for VOMCs while less so for FOMCs. As indicated in Theorem 2, the information of counting statistics $\vec{n}_{i,s_{i,l}}$ is important since the weights heavily depend on $\vec{n}_{i,s}(a)$. Yet due to the softmax function, only (normalized) probability values can be obtained instead of the counts. With the backward statistics $\vec{g}^{\leftarrow}_{i,s}$, $\vec{n}_{i,s_{i,l}}$ can be derived as
@@
Taking $M=M^{\prime}-1=D$, after the statistics collection layer, a sufficiently wide $F$ layer with ReLU activation can compute (by the universal approximation capability) (Cybenko,, 1989; Hornik et al.,, 1989).
@@
There exists an AA-head transformer
layer that can perform the induction: Takes $\vec{H}^{(\ell)}$ in (7) as input and outputs $\vec{H}^{(\ell+1)}$, and the final output layer taking $\vec{H}^{(D+3)}$ as input can output the AA-dim Bayesian optimal vector $\sum_{l=0,\ldots,D}~\omega_{n,l}\vec{p}_{n,s_{n,l}}.$
@@
The construction is certainly not unique. Consider a 2-layer transformer with a large embedding space, where the number of attention heads scales like $|\mathcal{A}|^{D}$. Each attention head could be precoded to collect the statistics for each possible suffix of length $\leq D$, and all such information can be stored in $\vec{h}_{i}^{(3)}$, assuming a sufficiently large embedding space (or a more delicate embedding structure in a smaller space). Each $\vec{h}_{i}^{(3)}$ would then have all the information on the context tree, on which a wide $F$ layer could perform the full CTW algorithm to produce the next token distribution. We conscientiously forgo constructions like this, because in large models with $\mathcal{A}$ in the tens of thousands and $D$ in the tens or hundreds, scaling as $|\mathcal{A}|^{D}$ is unrealistic, and the computation of the $F$ layer would be extremely complex and hard to interpret. More discussions on potential tradeoffs between layers and attention heads can be found in Rajaraman et al., 2024a .
@@ 4.5 Interpreting Two-Layer Transformers
We next consider how two-layer transformers can learn VOMCs in context. Our construction given above requires $D+2$ layers, and we conjecture that the statistical information collected in the second layer can be used to approximate the optimal weights $\omega_{n,l}$ with an $F$ network.
To verify our conjecture, we implement the first two constructed transformer layers but control the statistics passed onto the $F$ layer. In Fig. 8, “TF 0-2” denotes the canonical 2-layer transformer; “TF 2-0” is the transformer with two constructed layers with output $\vec{a}_{i}^{(2)}$ in (6), followed by a trainable $F$ layer (the $F$ layer in the second layer of the transformer) and an output layer. The version “TF 2-0 w/o counts” does not contain $\vec{g}_{i-1,M^{\prime}}^{\leftarrow}$ or $\vec{pos}_{i}$ in $\vec{a}_{i}^{(2)}$; the version “TF 2-0 total counts only” does not contain $\vec{g}_{i-1,M^{\prime}}^{\leftarrow}$ in $\vec{a}_{i}^{(2)}$ and $\vec{pos}_{i}$ is replaced by the total count $i$; “TF 2-0 w/ all counts” replaces $\vec{g}_{i-1,M^{\prime}}^{\leftarrow}$ and $\vec{pos}_{i}$ with $\{\vec{n}_{n,s_{n,l}}\}_{l=0}^{D}$ and ii. Even though their performances are rather clustered, we can make the following observations: 1) The performances degrade as more counting information is removed from the representation, and the counting information is clearly important, 2) The performances of “TF 2-0” and “TF 2-0 w/ all counts” almost match exactly, indicating the main function of the backward statistics $\vec{g}^{\leftarrow}_{i-1,M^{\prime}}$ is to extract the counts, and 3) The performance of the canonical 2-layer transformer is similar to that of the constructed “TF 2-0” and “TF 2-0: w/ all counts” than those with less counting information. Therefore, the performances of these different implementations confirm the conjecture. Interestingly, Akyürek et al., (2024) proposed a similar learned-weight 22&33-gram algorithm to interpret how transformers learn finite automata, and suggested it may be optimal. In contrast, we obtain the (suboptimal) architecture by approximating the provable optimal transformer architecture.
@@ Appendix D Training Details
We choose the alphabet size to be $|\mathcal{A}|=3$ in the experiments. For training, we randomly generate $K=20000$ CTs of various depths (maximum order $D\leq 5$), and then for each CT leaf, we generate a probability distribution. Two different ways of generating these probability distributions are taken: the first approach is to use the Dirichlet distribution to sample such distributions, and the second approach is to randomly select some of the elements in the alphabet to have probability zero, and the others with i.i.d. random values before normalization. Different values of the Dirichlet parameter are tested but only the results do not appear to be sensitive to the choice. For each CT, a source sequence of a certain length (e.g., $N_{k}=5120$) is produced. The context window NN can vary, but in most cases, we set it at $512$ (except when $D=5$, we set it to be $1536$ to allow sufficient data collection in context). Each source sequence is segmented into $\lfloor N_{k}/N\rfloor$ training sequences.
@@ G.1 A New Formula for Bayesian Next Token Prediction
Discussion. Note that $p^{e}_{n,s},p^{w}_{n,s}$ can be efficiently calculated by the CTW procedure, and compared to calculate $\frac{P_{\pi_{\text{CTW}}}(x_{1}^{n+1}|x_{1-D}^{0})}{P_{\pi_{\text{CTW}}}(x_{1}^{n}|x_{1-D}^{0})}$ for each $x_{n+1}$ the extra computation besides the CTW procedure is AA times larger than that by Eq (10). The weighted average formula in Eq (10) gives a natural interpretation for the Bayesian optimal next token predicted probability. Each suffix along the root the leaf path $s_{n,0}-s_{n,1}-\cdots-s_{n,D}$ can potentially be the true suffix, i.e., $s_{n,l}\in\mathcal{L}(T)$, and $\vec{p}_{n,s_{n,l}}$ is in fact the Bayesian optimal next token prediction given $s_{n,l}\in\mathcal{L}(T)$.
The blending weights $\omega_{n,l}$’s are based on stopping probability $\lambda$, the information in the potential suffix path such as $p^{e}_{s_{n,s_{n,l}}}$ as well as the information from their siblings $p^{w}_{n,qs_{n,l-1}}$. We can interpret $p^{e}_{n,s}$ as the evidence (unnormalized likelihood) that $s\in\mathcal{L}(T)$, and $p^{w}_{n,s}$ as the evidence that $s\in T$, i.e., the underlying tree covers node $s$. Theorem 6 indicates that more weights are assigned to $s_{n,l}$ than $s_{n,l-1}$, i.e., $\omega_{n,l}>\omega_{n,l-1}$, if $\lambda$ is smaller (i.e., node $s_{n,l-1}$ is more likely to branch and thus less likely to be a leaf node), $p^{e}_{n,s_{n,l}}-p^{e}_{n,s_{n,l-1}}$ is larger (i.e., $s_{n,l}$ has more evidence than $s_{n,l-1}$) and $\sum_{q\in\mathcal{A}}\ell^{w}_{n,qs_{n,l-1}}-\ell^{w}_{n,s_{n,l}}$ is larger (i.e., $_{n,l}’s s$iblings have more evidence to explain the data and thus $s_{n,l-1}$ is less likely to be a leaf node). The indicator function $\mathbb{I}_{l=D}$ is due to the maximum depth constraint on the branching process. Nodes at level $l=D$ automatically stop the branching, i.e., the branching-stopping probability is 11 for such nodes.